Find the subgame perfect nash equilibrium (SPNE) of the problem:

2 firms are competing (Stackelberg): 1st firm 1 picks q1, then firm 2 observes q1 and picks q2. Price is determined by P(Q) = 40 - q1 - q2.

Firms 1 and 2 both have a fixed or marginal cost of 8.
Answer #1

Ans! P = 40-91-92 profit of firm 2 T = (40-91-92) 92 - 892 T2 = 4092-9, 92-92² – 892 F.OC of fit maximizatren, d 2 = 40 - q -